OpenStudy (briensmarandache):

can someone with trig experience check my work for sine2x cosine x + cosine2x sine x = .5sqrt(3)

OpenStudy (briensmarandache):

posting now

OpenStudy (freckles):

hint sin(a+b)=sin(a)cos(b)+cos(a)sin(b)

OpenStudy (freckles):

oh ok

OpenStudy (briensmarandache):

one more page

OpenStudy (freckles):

\[\sin(2x+x)=\frac{\sqrt{3}}{2}\] \[\sin(3x)=\frac{\sqrt{3}}{2}\] so this is the equation you solved?

OpenStudy (freckles):

oh mine it looks like you took a more complicated route

OpenStudy (briensmarandache):

is it incorrect? what is the easier process? lol

OpenStudy (freckles):

what I did above use the sum identity for sin

OpenStudy (freckles):

but I will look at yours

OpenStudy (freckles):

if you have a*b=0 then you can do the whole a=0 or b=0

OpenStudy (freckles):

you can't do if a*b=sqrt(3)/2 then one of the factors has to be sqrt(3)/2 because that isn't necessarily true

OpenStudy (freckles):

do you know how to solve sin(3x)=.5*sqrt(3) ?

OpenStudy (freckles):

you would first find what u satisfies the equation sin(u)=.5*sqrt(3) by using unit circle

OpenStudy (freckles):

then you would divide answer by 3 to get what x is since u=3x

OpenStudy (briensmarandache):

ok, let me try that

OpenStudy (briensmarandache):

pi/6

OpenStudy (briensmarandache):

do you want to post my work?

OpenStudy (briensmarandache):

me*

OpenStudy (freckles):

you are looking for when the y-coordinate is sqrt(3)/2

OpenStudy (freckles):

for what angles is the y-coordinate sqrt(3)/2

OpenStudy (briensmarandache):

so x is not equal to pi/6?

OpenStudy (briensmarandache):

let me post my work so u can see what i did

OpenStudy (freckles):

alright

OpenStudy (briensmarandache):

posting now

OpenStudy (briensmarandache):

one min this site is not very well developed

OpenStudy (freckles):

yeah it has been pretty slow

OpenStudy (briensmarandache):

i have to upload it to my gdrive and then post it

OpenStudy (briensmarandache):

but do you get pi/6 for the answer

OpenStudy (freckles):

the y-coordinate is sqrt(3)/2 when you have the angles pi/3 or 2pi/3 and you can do +2npi if you wanted

OpenStudy (freckles):

\[u=\frac{\pi}{3}+2n \pi \text{ or } u=\frac{2\pi}{3}+2n \pi\] but u=3x

OpenStudy (freckles):

so divide both sides by 3

OpenStudy (freckles):

3*3=9 not 6

OpenStudy (briensmarandache):

lol wow ok

OpenStudy (freckles):

\[\frac{\pi}{3} \cdot \frac{1}{3}=\frac{\pi \cdot 1}{3 \cdot 3}=\frac{\pi}{9}\]

OpenStudy (freckles):

but anyways that is just one value that satisfies the equation

OpenStudy (freckles):

there are infinitely more solutions to this equation

OpenStudy (briensmarandache):

so other that that, its good? pi/9 is one solution an then i do the same thing for sin 2x+x?

OpenStudy (freckles):

sin(3x)=sqrt(3)/2 gives us two solutions for 3x in the first rotation of a circle 3x=pi/3 or 3x=2pi/3 correct?

OpenStudy (briensmarandache):

so the answer would be infinity many solutions, he asks for all real solutions.

OpenStudy (freckles):

first of all do you understand my solutions for 3x the first rotation ?

OpenStudy (briensmarandache):

yes

OpenStudy (freckles):

and then for any more rotations you can do +2 pi*n 2*pi carries us once around the circle 2*pi*2 carries us twice around the circle 2*pi*3 carries us three times around the circle ... 2*pi*n carries us n times around the circle where n is an integer also sin and cos have period 2pi ---- so anyways we have 3x=pi/3+2pi*n or 3x=2pi/3+2pi*n

OpenStudy (freckles):

and then to solve for x you divide both sides by 3

OpenStudy (freckles):

\[x=\frac{\pi}{9}+\frac{2 \pi n }{3} \\ x=\frac{2 \pi}{9}+\frac{2 \pi n}{3} \\ \text{ where n is an integer }\]

OpenStudy (freckles):

that gives you all the solutions

OpenStudy (freckles):

for sin(2x)cos(x)+sin(x)cos(2x)=sqrt(3)/2

OpenStudy (freckles):

@briensmarandache do you have any questions did i like totally make you speechless

OpenStudy (freckles):

or typeless in this case

OpenStudy (briensmarandache):

i feel like this is a stupid question so i was sitting on it for a bit

OpenStudy (freckles):

I'm sure whatever question you have isn't stupid

OpenStudy (briensmarandache):

so if i did the same thing for sine 2x+x = sqrt(3)/2, i would get the same solution for x or a solution that looks different but means the same thing or something different all together?

OpenStudy (freckles):

for sin(2x+x)=sqrt(3)/2?

OpenStudy (freckles):

2x+x is 3x

OpenStudy (freckles):

sin(2x+x)=sqrt(3)/2 is the same equation as sin(3x)=sqrt(3)/2

OpenStudy (freckles):

and both of those are the same equation as sin(2x)cos(x)+cos(2x)sin(x)=sqrt(3)/2

OpenStudy (briensmarandache):

i see now, ok. lol, it was a stupid question! thanks!!!

OpenStudy (freckles):

sin(a+b)=sin(a)cos(b)+cos(a)sin(b) so sin(a)cos(b)+cos(a)sin(b)=sin(a+b) sin(2x)cos(x)+cos(2x)sin(x)=sin(2x+x)=sin(3x) I used that a was 2x and b was x

OpenStudy (freckles):

So first step to solve this equation you have here is to realize you can use the sum identity for sin

OpenStudy (freckles):

what if I had this: \[\cos(2x)\cos(x)-\sin(2x)\sin(x)=\frac{\sqrt{3}}{2}\]

OpenStudy (briensmarandache):

that would be cos(x+y)

OpenStudy (freckles):

right and you would have this as a result: \[\cos(2x+x)=\frac{\sqrt{3}}{2} \\ \cos(3x)=\frac{\sqrt{3}}{2} \\ 3x=\frac{\pi}{6}+2n \pi \\ 3x=\frac{-\pi}{6}+2n \pi\] where n is integer

OpenStudy (freckles):

or instead of -pi/6 you could have said 11pi/6 which ever one i like to use cos is even when I can tht is cos(x)=cos(-x)

OpenStudy (freckles):

anyways I'm glad you understand more

OpenStudy (briensmarandache):

thanks this really helped a lot!!!!

OpenStudy (freckles):

have fun and trig should be a blast !

OpenStudy (briensmarandache):

i will, i like it already. once i have a better understanding of everything im sure ill love it more lol